### v1

# Benchmarking Large Language Models for Math Reasoning Tasks

Abstract
The use of Large Language Models (LLMs) in mathematical reasoning has become a cornerstone of related research, demonstrating the intelligence of these models and enabling potential practical applications through their advanced performance, such as in educational settings. Despite the variety of datasets and in-context learning algorithms designed to improve the ability of LLMs to automate mathematical problem solving, the lack of comprehensive benchmarking across different datasets makes it complicated to select an appropriate model for specific tasks. In this project, we present a benchmark that fairly compares seven state-of-the-art in-context learning algorithms for mathematical problem solving across five widely used mathematical datasets on four powerful foundation models. Furthermore, we explore the trade-off between efficiency and performance, highlighting the practical applications of LLMs for mathematical reasoning. Our results indicate that larger foundation models like GPT-4o and LLaMA 3-70B can solve mathematical reasoning independently from the concrete prompting strategy, while for smaller models the in-context learning approach significantly influences the performance. Moreover, the optimal prompt depends on the chosen foundation model. We open-source our benchmark code to support the integration of additional models in future research.
Introduction
Increasing computational resources and the free accessibility to large amounts of textual data have boosted the developments of Large Language Models (LLMs) (Brown et al. 2020; OpenAI 2023a; Touvron et al. 2023; Gemini-Team et al. 2023) in recent years. Central to this progress is the transformer architecture (Vaswani et al. 2017) based on the self-attention mechanism, which superseded recurrent neural networks for natural language processing and led to a trend of extremely large and powerful models, mastering many text-based tasks, such as summarizing (Zhang et al. 2024; Goyal, Li, and Durrett 2022), translating (Jiao et al. 2023) or question answering (Chowdhery et al. 2023).
While these models perform remarkably well for many language tasks, solving mathematical word problems still presents a challenge for LLMs (Hendrycks et al. 2021; Imani, Du, and Shrivastava 2023; Kaddour et al. 2023). Such problems require not only a deep understanding of the task description but also sophisticated logical and mathematical reasoning to determine the solution path. Additionally, the computational demands of finding the result pose further challenges. Since foundational models are not specifically optimized for mathematical calculations, they often struggle even with elementary arithmetic (Cobbe et al. 2021). Figure 1 lists various answers from different models.
The ability to perform mathematical reasoning is crucial not only in educational settings, where LLMs could potentially serve as a technological foundations for personalized tutoring (Liang et al. 2023; Kasneci et al. 2023), but also in higher education, where it could assist with complex problem-solving and discovering new proofs (Wang et al. 2024). Moreover, in professional fields such as finance and economics, where mathematical problems are prevalent, these capabilities could have significant applications (Korinek 2023).Therefore, enhancing LLMs’ mathematical reasoning skills would benefit various applications, extending from classroom support to advanced research and industry.
To address this challenge, several algorithms and methods have been proposed to improve the mathematical reasoning capabilities of LLMs. Understanding the strengths and limitations of each algorithm is essential for making informed decisions and effectively applying them in different contexts. While previous research has often evaluated these methods on rather limited datasets (Qiao et al. 2023) or using similar methods (Luo et al. 2023), there remains a need for more comprehensive analysis. Addressing this research gap, our study presents an extensive benchmarking of seven advanced methods across five widely-used datasets. Additionally, whereas prior work has mainly focused on accuracy as the main performance metric, we also consider time and cost factors, for deploying these algorithms effectively.
More specifically, to tackle mathematical reasoning with LLMs, we identify three core tasks: First, achieving high accuracy requires the model to correctly and consistently explain the reasoning path leading to the correct result. Second, robustness is demonstrated through the model’s ability to produce the correct result across multiple repetitive calls. Finally, efficient resource usage - encompassing time and API costs - is crucial for practical deployment.
To tackle these challenges, there are different ideas for improving the mathematical reasoning capabilities of LLMs. Keeping the resource consumption low, the prompt for the LLM can be adapted by providing various samples and initiating a step-by-step problem solving approach. The Chain-of-Thought (CoT) approach (Wei et al. 2022) as well as the Auto CoT (Zhang et al. 2023) exploit this idea. In a similar way, the Zero-Shot CoT (Kojima et al. 2022) operates, but without the few-shot samples. Setting the focus on the robustness of the solution and therefore requiring a higher computational effort, the Self-Consistency (Wang et al. 2023) and Complex CoT (Fu et al. 2023) approaches optimize the solving process. Both call the foundational model repeatedly and therefore increase the robustness. Lastly, external engines like Python can be exploited outsourcing the computational tasks partly from the LLM to an external resource. PAL (Gao et al. 2023) and PoT (Chen et al. 2023) are both using this approach, applying few-shot samples and pipelines adapted for this use case. All these approaches bring their trade-off between accuracy, robustness, and consumption of resources, and for a fair comparison, all dimensions need to be investigated thoroughly.
We aim to bridge this gap by comparing different mathematical reasoning strategies based on LLMs in multiple dimensions, providing a trade-off between costs and performance. We make our source code publicly available to enable research to reproduce our findings. In summary, our contributions are as follows:
We provide a detailed empirical evaluation of seven approaches in multiple dimensions, providing a trade-off between performance, robustness and consumption of resources (time and costs).
Our evaluation results highlight several merits, such as identifying that the algorithms Auto CoT, together with LLaMA 3-70B, and Zero-Shot CoT with GPT-3.5, provide the optimal trade-off between performance and computational resources.
We release an open-source benchmarking codebase for fair mathematical reasoning, which can be easily extended for future research.
Related Work
Multiple surveys examine the problem of mathematical reasoning with LLMs from theoretical viewpoints. Lu et al. (2023) investigate the different deep learning methods used for mathematical reasoning, Ahn et al. (2024) give a thorough overview specifically covering LLMs for this task, and Huang and Chang (2023) analyse the different kinds of reasoning problems and the corresponding opportunities for LLMs. All surveys emphasize the missing generalizability and robustness, the challenges with complex questions, and the problem of hallucinations and trustworthiness.
Setting the focus on the algebraic skills of LLMs, Yuan et al. (2023) find that GPT-3.5 and GPT-4 outperform the other foundation models. As key points, they mark the influence of tokenization, the importance of the model size, and the sensitivity of prompts, identifying different optimal prompts across the LLMs. However, they also emphasize that the algebraic ability of a model cannot be directly equated with the mathematical reasoning capabilities.
From an empirical viewpoint Qiao et al. (2023) show that Chain-of-Thought reasoning works decently on the GSM8K dataset for GPT-3 (Brown et al. 2020) and Codex (Chen et al. 2021), but poor for LaMDA (Thoppilan et al. 2022), and the results for PaLM (Chowdhery et al. 2023) lie in-between. They compare different model sizes, in general, finding that bigger models perform better. Luo et al. (2023) compare various foundation models on the GSM8K and the MATH dataset, finding that closed-source models perform significantly better than open-source models and GPT-4 being the overall best model for mathematical reasoning.
A detailed analysis of various LLMs for mathematical reasoning tasks in combination with visual inputs is reported by Lu et al. (2024). In line with the previous results, they show the best performance for GPT-4 Vision (OpenAI 2023b), which lies still clearly below human achievements.
All of the studies do not report any reference time or resource consumption of the different models to compare them from this point of view. We aim to close this gap by comparing different mathematical reasoning methods empirically on various foundation models, not only comparing the performance but also reporting thoroughly the associated costs and time factors. Also, they compare the foundation models but not the different prompting strategies applied.
Benchmarking Details
The goal of this benchmarking is to systematically evaluate different strategies for fitting and aligning large language models (LLMs) for mathematical reasoning. We perform a comparative analysis of these methods across five datasets, using both open-source and closed-source ground models, and evaluate performance across multiple dimensions. The following sections provide a detailed description of our benchmarking methodology.
Datasets for Mathematical Reasoning
To assess the reasoning ability of the methods, we use several mathematical word problems. These tasks stem from five publicly available datasets that are commonly used in the relevant literature and represent varying levels of difficulty, from elementary school to university mathematics. The datasets also include various types of problems, including both open-ended questions and multiple-choice tasks. The specific details of the datasets are presented in Table 1.
Dataset | #Train | #Test | Problem Type | Level |
GSM8K | 7,470 | 1,319 | Word Problem | GS |
SVAMP | 700 | 300 | Word Problem | GS |
Multi-Arith | 420 | 180 | Word Problem | GS |
MATH | 7,500 | 5,000 | Word Problem | HS |
AQuA | 97,500 | 254 | Multiple Choice | CO |
Foundation Models
Many of the mathematical reasoning methods are model-agnostic, i.e. they are not dependent on a particular base model. Although most studies report results using GPT-3.5 and GPT-4, the choice of additional models varies. To ensure a fair and comprehensive comparison, we evaluate these methods using state-of-the-art open-source and closed-source base models. In addition to GPT-3.5 111gpt-3.5-turbo-0125 and GPT-4o 222gpt-4o-2024-05-13 (OpenAI 2023a), we analyze LLaMA 3-8B and LLaMA 3-70B models (Touvron et al. 2023).
Methods for Mathematical Reasoning
We compare in-context methods that keep the parameters of the foundation models unchanged while employing prompting strategies to enhance the models’ capabilities. These methods can be categorized into classic prompt engineering, process optimization, and the use of external engines.
The standard Chain-of-Thought (CoT) approach (Wei et al. 2022) involves providing the model with multiple examples where the solution process is detailed step by step, guiding the model to generate a reasoning chain before arriving at a conclusion. The Zero-Shot CoT approach (Kojima et al. 2022) streamlines this by replacing explicit examples with the key phrase "Let’s think step by step.". This strategy involves two model calls: one to generate the reasoning path and another to extract the final answer. A variant of the classic CoT is the Auto CoT approach (Zhang et al. 2023), which replaces manually curated few-shot examples with automatically selected ones. This is achieved by clustering the training data and selecting the best representative from each cluster.
Due to the stochastic nature of foundation models, the same prompt can yield both correct and incorrect solutions. To address this variability, the Self-Consistency CoT approach (Wang et al. 2023) involves generating multiple responses from the model and using a majority vote to select the most likely correct solution. This method utilizes the same few-shot examples as the standard CoT. In contrast, the Complex CoT approach (Fu et al. 2023) operates on the assumption that higher complexity can enhance performance. It selects the most complex examples from the training data for few-shot input and generates multiple solutions, applying majority voting only to the most complex responses.
Recognizing the limitations of LLMs in algebraic tasks, the Program-of-Thought (PoT) (Chen et al. 2023) approach extends CoT to make use of an external Python engine. In the few-shot examples, it is guided to separate the reasoning process from the computation part by generating program code. A similar idea is exploited by PAL (Gao et al. 2023), which integrates program synthesis into the model’s workflow, focusing on generating executable code to solve tasks.
GSM8K | SVAMP | Multi-Arith | AQuA | MATH | ||
CoT | GPT-3.5 | 0.78 0.34 | 0.86 0.30 | 0.99 0.09 | 0.74 0.34 | 0.57 0.45 |
GPT-4o | 0.97 0.18 | 0.96 0.18 | 1.00 0.00 | 0.87 0.30 | 0.79 0.38 | |
LLaMA 3-8B | 0.63 0.38 | 0.78 0.32 | 0.98 0.05 | 0.61 0.30 | 0.32 0.41 | |
LLaMA 3-70B | 0.87 0.29 | 0.91 0.23 | 1.00 0.01 | 0.60 0.41 | 0.41 0.43 | |
Auto CoT | GPT-3.5 | 0.71 0.35 | 0.90 0.27 | 0.92 0.22 | 0.76 0.35 | 0.53 0.45 |
GPT-4o | 0.97 0.18 | 0.96 0.17 | 1.00 0.00 | 0.86 0.30 | 0.80 0.37 | |
LLaMA 3-8B | 0.67 0.36 | 0.85 0.28 | 1.00 0.02 | 0.62 0.31 | 0.19 0.28 | |
LLaMA 3-70B | 0.92 0.22 | 0.93 0.21 | 1.00 0.00 | 0.79 0.33 | 0.49 0.45 | |
Zero-Shot CoT | GPT-3.5 | 0.91 0.23 | 0.91 0.25 | 1.00 0.03 | 0.80 0.30 | 0.48 0.46 |
GPT-4o | 0.93 0.25 | 0.93 0.24 | 1.00 0.01 | 0.81 0.36 | 0.61 0.46 | |
LLaMA 3-8B | 0.57 0.36 | 0.75 0.30 | 0.87 0.19 | 0.25 0.26 | 0.14 0.26 | |
LLaMA 3-70B | 0.81 0.33 | 0.85 0.25 | 0.98 0.06 | 0.26 0.34 | 0.19 0.33 | |
Complex CoT | GPT-3.5 | 0.90 0.26 | 0.86 0.31 | 1.00 0.03 | 0.51 0.41 | 0.54 0.45 |
GPT-4o | 0.96 0.20 | 0.95 0.20 | 1.00 0.00 | 0.80 0.37 | 0.63 0.46 | |
LLaMA 3-8B | 0.58 0.43 | 0.80 0.33 | 0.99 0.03 | 0.68 0.30 | 0.27 0.38 | |
LLaMA 3-70B | 0.92 0.28 | 0.91 0.26 | 1.00 0.00 | 0.83 0.30 | 0.47 0.45 | |
Self-Consistency CoT | GPT-3.5 | 0.65 0.39 | 0.80 0.39 | 0.97 0.15 | 0.71 0.43 | 0.27 0.41 |
GPT-4o | 0.96 0.20 | 0.95 0.20 | 1.00 0.00 | 0.87 0.32 | 0.81 0.38 | |
LLaMA 3-8B | 0.60 0.38 | 0.83 0.31 | 1.00 0.00 | 0.69 0.33 | 0.28 0.41 | |
LLaMA 3-70B | 0.88 0.32 | 0.94 0.22 | 1.00 0.00 | 0.65 0.43 | 0.44 0.45 | |
PAL | GPT-3.5 | 0.77 0.38 | 0.80 0.38 | 0.93 0.26 | 0.40 0.44 | N/a |
GPT-4o | 0.96 0.19 | 0.96 0.19 | 1.00 0.01 | 0.52 0.47 | N/a | |
LLaMA 3-8B | 0.57 0.40 | 0.74 0.36 | 0.91 0.22 | 0.24 0.35 | N/a | |
LLaMA 3-70B | 0.82 0.34 | 0.85 0.30 | 0.95 0.18 | 0.35 0.44 | N/a | |
PoT | GPT-3.5 | 0.75 0.44 | 0.87 0.33 | 0.99 0.10 | 0.50 0.50 | N/a |
GPT-4o | 0.83 0.38 | 0.80 0.40 | 1.00 0.00 | 0.66 0.47 | N/a | |
LLaMA 3-8B | 0.54 0.50 | 0.58 0.49 | 0.42 0.49 | 0.31 0.46 | N/a | |
LLaMA 3-70B | 0.82 0.38 | 0.87 0.33 | 0.70 0.46 | 0.58 0.49 | N/a |
Many methods can be combined or varied by adjusting the number of contextual examples, applying techniques such as majority voting, or incorporating additional post-processing steps to improve the results. To obtain a focused and practicable benchmark, we concentrate on the core concepts without implementing these variations. Figure 2 gives an overview of the different methods, highlighting the primary ideas and the main differences in the technical implementation, e.g. the use of examples in the prompts or the need for iterative optimization. It is important to note that iterative optimization requires more computational resources.
Performance Metrics
To evaluate the results, we use the pass@k metric (Chen et al. 2021), which measures the probability that the correct answer is generated at least once when the model is prompted times. Chen et al. (2021) introduced an unbiased estimator for this metric by prompting the model times, which helps mitigate the impact of the inherent stochasticity in language models. Given as the number of correct responses, the metric is calculated as follows:
When using , we report the averaged accuracy of the algorithms on solving mathematical problems. With a larger , we can study the robustness of different methods. Beyond performance in successfully solving the task, we also compare the resource consumption. For open-source models (LLaMA), we track elapsed time, while for closed-source models (GPT), we monitor the costs of API calls.
Experimental Results
To compare the different methods on the various foundational models, we conduct an extensive analysis over five datasets. In the following we describe the outcomes of our study. The code for our experiments can be found here 333https://github.com/kathrinse/Math-Reasoning-Benchmark.
Experimental setup
We integrated the original implementations of the papers into our repository if code is provided by the paper. Similarly, we took all the original few-shot examples as far as the papers provide. For Auto CoT and Complex CoT, we followed their instructions to generate the few-shot samples accordingly, and for both we manually added the missing reasoning paths for the samples. The Self-Consistency approach always used the same prompt as the vanilla CoT.
For the smaller models, we generated results per sample to compute the pass@3. The larger models (GPT-4o and LLaMA 3 - 70B) were stable enough to decrease for cost and resource savings. For Complex CoT, we generated 5 solutions and applied the majority voting on the 3 most complex ones. For the Self-Consistency CoT approach, we always took the majority of 10 repeated runs.
The questions in the MATH dataset range from algebraic up to geometric problems, often requiring more complex results than mere floating point numbers. Therefore, the methods relying on external engines to perform the main calculus, were not suitable for this advanced kind of problems.
The LLaMA models were executed locally on one (for 8B) resp. two (for 70B) NVIDIA A100 80GB GPUs, both ran in half-precision (float16). The batch size was equal to , except for Self-Consistency and Complex CoT, where it was equal to the number of runs for the majority voting. Concrete prompts and more details can be found in Appendix.
Robust Performance
In this section, we examine the influence of various prompt strategies, foundation models, and datasets on the pass@3 metric, which represents robust performance. Table 2 lists the comparison results. Besides the pass@3 metric, we also report the accuracy reached by all models in Appendix.
Foundation Models.
When comparing different foundation models, we see that algorithms incorporating GPT-4o generally achieve the best performance. Moreover, LLaMA 3-70B and GPT-3.5 are always in the second and third places, respectively, highlighting the connection between model capability and size. It is worth noting that the strong performance of LLaMA 3-70B is even more robust, while the outcomes of GPT-3.5 rely more on the quality of the prompt. For instance, on the dataset GSM8K, LLaMA 3-70B consistently achieves an accuracy of above 81%, while GPT-3.5 ranges from 65% (Self-Consistency CoT) to 91% (Zero-Shot CoT). A smaller model, LLaMA 3-8B, also fluctuates more depending on the concrete method applied.
Prompt Strategies.
Different foundation models favor different prompt strategies. As previously mentioned, GPT-4o combined with CoT outperforms other algorithms, while for both LLaMA 3 models, the Auto CoT approach yields the largest performance boost. For GPT-3.5, Zero-Shot CoT demonstrates the best performance, suggesting that additional shots may not benefit the model but need clarification. Instead, Zero-Shot CoT calls the foundation model twice per sample—first for reasoning and second for the final prediction—which appears to increase the stability of the final result. In contrast, all other models call the foundation models only once for reasoning. Despite this, the costs and time consumption of Zero-Shot CoT remain comparable due to its concise prompt. Additionally, the two methods using the external engine are less favorable in performance than other CoT-based methods, especially on the AQuA datasets. Here, the computation executed by the external engine is not optimal for the advanced reasoning required and the multiple choice structure of this dataset poses an additional challenge.
Datasets.
The Multi-Arith dataset raises less difficulties; even a small LLaMA 3 model exceeds 80% pass@3, while larger models can achieve near-perfect performance. Therefore, the differences between prompting strategies are negligible. The only exception is PoT, which, in its original code base, implements a zero-shot code generation approach that performs poorly across all foundation models.
In contrast, the AQuA dataset, due to its college-level difficulty, presents a greater challenge. Its multiple-choice format causes problems for code-generated solutions, making them unsuitable for this dataset. Here, the choice of prompting strategies significantly impacts the largest LLaMA 3 model, with the Zero-Shot CoT approach producing inferior results compared to GPT-4 using Complex CoT. GPT-3.5 also shows more variability in its performance.
In general, the quality of the foundation model has a more significant impact on performance. GPT-4o and LLaMA 3-70B consistently achieve over 80%, regardless of prompting strategy, on all datasets except for MATH. Here, even GPT-4o ranges from merely 61% up to 81%. LLaMA 3-70B achieves only 21% pass@3 applying Zero-Shot CoT and performs best employing the Auto CoT strategy with 49%. There is a significant performance difference among all foundation models, except between LLaMA 3-70B and GPT-3.5. For a detailed ANOVA analysis, see Appendix.
Efficiency
Beyond performance, computational cost is an important factor to consider in practice. In this section, we examine the trade-off between performance and required resources. Figure 3 illustrates a complete comparison of different prompt strategies on GSM8K. From the results on LLaMA 3, we see that larger model 70B achieves much higher result in pass@3, but also costs more time. For instance, Auto CoT with LLaMA 3-70B has a computation time of seconds per sample, while LLaMA 3-8B takes only second per sample. Comparing different prompt strategies, Auto CoT has the best trade-off between performance and costs, whereas Self-Consistency CoT is the most computationally costly method due to its multiple refinement runs.
When using GPT as our foundation model, the computational resources are measured by the costs in US Dollars for each sample. Similar to the findings from LLaMA 3 models, a larger model (GPT-4o) performs better than a smaller one (GPT-3.5) but is more computationally costly. However, GPT-3.5 with Zero-Shot CoT achieves a good trade-off, as its performance reaches almost the same level as GPT-4o on GSM8K, but its costs are approximately 30 times less. Auto CoT with GPT-4o increases the accuracy of pass@3 by four percentages compared to GPT-3.5 with Zero-Shot CoT, but it also requires approximately 20 times more US Dollars. Therefore, GPT-3.5 with Zero-Shot CoT obtains the best trade-off, highlighting its potential for practical uses.
LLaMA 3 with Auto CoT consistently shows its advantages in high efficiency and performance on the other four datasets. GPT-3.5 with Zero-Shot CoT is also an optimal option except for the dataset MATH, where GPT-4o with Auto CoT reveals a significant performance improvement. Detailed illustrations can be found in Appendix.
Proficiency
The MATH dataset consists of questions from different domains (e.g. Algebra or Geometry) and levels (from 1 to 5). Figure 4 gives an comprehensive overview of the performance of the LLaMA 3 models using the CoT prompting strategy. According to this analysis the level 1 questions in Algebra and Prealgebra can be solved frequently, of the time by the all models. As also shown in Figure 4 the 70B model reaches an higher performance for all domains compared to 8B model. On the other hand, the level 4 and 5 questions pose a challenge for both models with less then 30% correct answers for nearly all domains. This differentiation shows the strength of the MATH dataset, allowing for a detailed analysis, and also highlight the need for further research in more complex and higher-level mathematical settings.
Figure 1 qualitatively compares the answers generated by all models for a question in the Algebra domain of level 1 and Figure 5 analyses a level 5 question for both LLaMA 3 models. The 8B model solved the level 1 task correctly but struggled with the more complex level 5 question. Concretely, the answers from the 8B and 70B models are the same for the level 1 question, and they are highly consistent with the true answer. However, the challenge for the 8B model to answer the question of level 5 is to correctly identify the first step, converting the question into an inequality. Another common failure of the small model is the limitation in calculations. In this case, LLaMA 3-8B wrongly calculates the value of for . The 70B model improves the ability to understand the question and mathematical calculation, leading to a correct answer.
Both GPT models yield in general a high proficiency over all types and levels. As shown in Appendix, GPT-3.5 reaches in level 5 between 11% and 52% pass@3, and GPT-4o even surpasses 30% across all question types. However, it also requires a greater amount of tokens for its comprehensive argumentation path, associated with higher costs.
Discussion
In our benchmarking study, we provide empirical evidence that the choice of the foundation model significantly influences the performance of LLMs on mathematical reasoning tasks, overshadowing the impact of specific prompting strategies. For instance, models like GPT-4o and LLaMA 3-70B show performance enhancements of up to 20% over smaller models across various datasets, highlighting the crucial role of model architecture and size. These findings align with previous results, indicating that model size significantly enhances the performance, and GPT-4 proving to be the strongest model (Qiao et al. 2023; Luo et al. 2023).
For elementary school tasks, GPT-4o and LLaMA 3-70B were particularly effective, suggesting that their advanced architectures are well suited to solving simpler math reasoning tasks. GPT-4o consistently answered over 95% of the questions in the GSM8K dataset correctly using almost all methods, albeit with a significant drop to 83% accuracy for the reasoning program method (PoT). This drop can often be attributed to the verbose output of GPT-4o, which exceeds typical token constraints and requires a higher token allowance for conclusive answers. This behavior underscores the need to optimize token efficiency in practical applications, especially in situations where computational resources are limited. An illustrative example of this token constraint and its impact is given in the Appendix.
The main limitations to the use of GPT-4o and LLaMA 3-70B are their high operating costs and runtimes, which make them less suitable for routine tasks. For cost-sensitive contexts, GPT-3.5 is a more cost-effective alternative, offering significant savings in computational overhead while providing competent performance for simpler tasks. Similarly, LLaMA 3-8B excels on less demanding tasks despite its smaller number of parameters, although its utility decreases with increasing task complexity due to stability issues.
Our analysis further shows that there is no consistently superior prompting strategy for all datasets and models. Extended reasoning strategies appear to be particularly beneficial for improving the performance of smaller models, while the differences between strategies for robust models such as GPT-4o and LLaMA 3-70B are relatively small. For the LLaMA models, Auto CoT provides an optimal balance between cost and performance, while Zero-Shot CoT has shown promising results on several tasks.
Despite these advances, there is still room for improvement in the performance of mathematical reasoning tasks at school and university level. For example, while GPT-4o achieves over 80% accuracy in the AQuA dataset, human performance is typically 85% according to (Zhong et al. 2023). In the MATH dataset, the maximum accuracy of GPT-4o is 75%, while PhD students achieve between 40% and 90% (Hendrycks et al. 2021). These results suggest that further progress in model training and reasoning skills is needed to close this gap, especially for tasks requiring deeper cognitive abilities and complex problem-solving.
These findings reveal important practical implications for real-world applications. In an educational context, GPT-3.5 can serve as a cost-effective solution for lower grades, while the more advanced GPT-4o model is better suited for high school mathematics. For data-sensitive applications, such as finance, LLaMA 3 offers a privacy-preserving and powerful open-source alternative. Furthermore, when implementing LLMs to solve these tasks, users might encounter issues such as occasional errors in reasoning and challenges in extracting relevant information from very long responses.
Conclusion
Our benchmarking study has shown that the choice of base model has a greater impact on the performance of LLMs on mathematical reasoning tasks than specific prompting strategies. Our results further suggest that while larger models such as GPT-4o and LLaMA 3-70B provide robust performance on simpler tasks, they present challenges in terms of scalability and efficiency, especially for complex problems like in higher education contexts. For cost-effective applications, GPT-3.5 and LLaMA 3-8B offer alternatives for less complex tasks, but show limitations for advanced problems.
Limitations.
We identify the limited number of foundation models used in our benchmark. Future work should consider other state-of-the-art open-source models such as Mistral-AI (Jiang et al. 2023) and closed-source models like Gemini (Gemini-Team et al. 2023) or Claude (Anthropic 2024). Nevertheless, our codebase has been released and is easy to use for integrating new foundation models. Moreover, we identified another set of models fine-tuned on mathematical reasoning problems, which we plan to include in our benchmarking for a more comprehensive evaluation beyond just prompt strategies.
Appendix A Appendix
Proficiency Analysis
To further analyze the proficiency of GPT models on the MATH dataset, we compare the results across different levels and question types in Figures 6 and 7. GPT-3.5 demonstrates relatively high performance in Algebra and Pre-algebra questions, but for the remaining categories, its pass@3 metric falls below 50% at Levels 4 and 5. In contrast, GPT-4 shows strong performance across all types and levels, with only Level 5 questions in Geometry, Intermediate Algebra, and Pre-calculus scoring below 50% pass@3.
This proficiency can be observed in the following example of a Level 5 question from the Algebra category of the MATH dataset. Both GPT models successfully solve the problem, though GPT-4o again demonstrates its talkative nature requiring many tokens to arrive at the final result.
Restricted number of tokens
Across all datasets and methods, the GPT-4 model struggles with some questions, not due to poor reasoning quality, but because it requires too many tokens for its argumentation. The following example from the GSM8K dataset demonstrates how the model’s detailed thinking leads to exceeding the token limit before arriving at an answer when using the PoT method. For comparison, we have also included the GPT-3.5 output.
Statistical analysis
To test our hypothesis that the differences between the foundation models are significant, while between the methods they are not, we conduct an ANOVA analysis on the results of the MATH dataset. The analysis for the foundation models yields a p-value of 1.9e-5, indicating a highly significant difference between them. In contrast, the analysis for the methods resulted in a p-value of 0.72, suggesting that the differences between methods are not statistically significant.
We also conduct a more detailed analysis of the differences between the foundation models. As illustrated in Figure 8, the results reveal that the difference between GPT-3.5 and LLaMA 3-70B is the only comparison that is not statistically significant. This suggests that while most foundation models show significant performance variations, GPT-3.5 and LLaMA 3-70B are more comparable in their results on the MATH dataset. This finding further emphasizes the importance of choosing the right foundational model, as it has a substantial impact on performance.
Accuracy results
Table 3 lists all accuracy results, corresponding to the pass@1 metric. The distribution of these results is similar to the pass@3 metric. GPT-4o with Self-Consistency CoT consistently outperforms all other methods and foundational models.
GSM8K | SVAMP | Multi-Arith | AQuA | MATH | ||
CoT | GPT-3.5 | 0.60 0.35 | 0.73 0.34 | 0.95 0.15 | 0.54 0.36 | 0.41 0.40 |
GPT-4o | 0.93 0.21 | 0.94 0.22 | 1.00 0.00 | 0.75 0.34 | 0.71 0.40 | |
LLaMA 3-8B | 0.39 0.31 | 0.56 0.33 | 0.83 0.17 | 0.31 0.23 | 0.14 0.23 | |
LLaMA 3-70B | 0.67 0.32 | 0.75 0.28 | 0.92 0.10 | 0.34 0.30 | 0.23 0.29 | |
Auto CoT | GPT-3.5 | 0.47 0.33 | 0.83 0.31 | 0.77 0.27 | 0.58 0.37 | 0.36 0.38 |
GPT-4o | 0.88 0.24 | 0.94 0.21 | 0.99 0.04 | 0.65 0.33 | 0.70 0.39 | |
LLaMA 3-8B | 0.43 0.32 | 0.67 0.33 | 0.90 0.13 | 0.32 0.23 | 0.08 0.14 | |
LLaMA 3-70B | 0.78 0.30 | 0.84 0.26 | 0.99 0.05 | 0.53 0.33 | 0.31 0.35 | |
ZeroShot CoT | GPT-3.5 | 0.82 0.29 | 0.82 0.30 | 0.97 0.10 | 0.60 0.35 | 0.34 0.39 |
GPT-4o | 0.79 0.31 | 0.91 0.26 | 0.97 0.08 | 0.63 0.36 | 0.50 0.43 | |
LLaMA 3-8B | 0.31 0.25 | 0.46 0.28 | 0.56 0.23 | 0.10 0.11 | 0.06 0.14 | |
LLaMA 3-70B | 0.55 0.32 | 0.59 0.29 | 0.80 0.17 | 0.10 0.16 | 0.08 0.18 | |
Complex CoT | GPT-3.5 | 0.81 0.31 | 0.78 0.35 | 0.97 0.10 | 0.32 0.32 | 0.38 0.40 |
GPT-4o | 0.89 0.25 | 0.94 0.22 | 0.80 0.00 | 0.59 0.31 | 0.55 0.46 | |
LLaMA 3-8B | 0.36 0.34 | 0.64 0.36 | 0.93 0.13 | 0.38 0.26 | 0.14 0.24 | |
LLaMA 3-70B | 0.74 0.32 | 0.82 0.31 | 0.99 0.04 | 0.59 0.34 | 0.31 0.37 | |
SelfConsistency CoT | GPT-3.5 | 0.34 0.22 | 0.77 0.40 | 0.96 0.17 | 0.66 0.44 | 0.19 0.33 |
GPT-4o | 0.95 0.21 | 0.95 0.21 | 1.00 0.00 | 0.81 0.35 | 0.76 0.39 | |
LLaMA 3-8B | 0.31 0.21 | 0.72 0.37 | 1.00 0.01 | 0.45 0.33 | 0.18 0.31 | |
LLaMA 3-70B | 0.82 0.35 | 0.89 0.26 | 1.00 0.01 | 0.46 0.38 | 0.30 0.37 | |
PAL | GPT-3.5 | 0.68 0.40 | 0.76 0.40 | 0.92 0.27 | 0.30 0.40 | N/a |
GPT-4o | 0.88 0.25 | 0.89 0.23 | 0.90 0.15 | 0.36 0.37 | N/a | |
LLaMA 3-8B | 0.36 0.33 | 0.55 0.36 | 0.70 0.26 | 0.13 0.23 | N/a | |
LLaMA 3-70B | 0.65 0.36 | 0.72 0.35 | 0.88 0.23 | 0.22 0.33 | N/a | |
PoT | GPT-3.5 | 0.75 0.44 | 0.87 0.33 | 0.99 0.10 | 0.50 0.50 | N/a |
GPT-4o | 0.83 0.38 | 0.80 0.40 | 1.00 0.00 | 0.66 0.47 | N/a | |
LLaMA 3-8B | 0.54 0.50 | 0.58 0.49 | 0.42 0.49 | 0.31 0.46 | N/a | |
LLaMA 3-70B | 0.82 0.38 | 0.87 0.33 | 0.70 0.46 | 0.58 0.49 | N/a |
Visualized Efficiency Comparison
To balance performance and resource consumption, we present the results across all datasets visually in Figures 9, 10, 11 and 12. On the smallest dataset, Multiarith, which is at the grade school level, nearly all methods and models consistently achieve over 90% pass@3. Even the smaller LLaMA 3-8B model proves to be an efficient and cost-effective option, particularly when used with the AutoCoT method. For the SVAMP dataset, the results are similarly strong, with pass@3 rates consistently exceeding 70%. Although GPT-4o and LLaMA 3-70B slightly outperform the other models, the margin is small.
In contrast, the results for the more advanced, college-level AQuA dataset are more varied. Interestingly, GPT-3.5 with Zero-Shot CoT nearly matches the best performance of GPT-4o while being significantly more resource-efficient. For LLaMA 3-70B, AutoCoT and Complex CoT deliver the best results.
The MATH dataset reveals the largest disparity between open-source and closed-source models. While LLaMA 3 struggles to surpass the 50% mark, GPT-4o consistently exceeds 60%, and even GPT-3.5 achieves between 50% and 60% pass@3.
Few-shot examples
For reproducibility, we include all few-shot examples used in the benchmarking survey, which are also available in our GitHub repository. Merely Zero-Shot CoT did not use any specific examples but only the prompt Let’s think step by step. For CoT, we used the prompts from the original paper (Wei et al. 2022), and the same prompts were used for Self-Consistency, see Figures 13 and 14. For MATH dataset we used the first 8 samples from the according training dataset.
As suggested by the original paper, we extracted the samples for the Auto CoT method using a clustering approach (Zhang et al. 2023). All used samples can be found in Figures 22, 23, 24, 25, 26 and 27. The Complex CoT approach extracts the “most complex” samples from the training data, to guide the reasoning process (Fu et al. 2023). We report our extracted samples in Figures 28, 29, 30, 31, 32 and 33. Due to limited GPU space, we used only 6 shots for the MATH dataset in the Complex CoT approach.
For PAL, we use the prompts suggested by the original paper (Gao et al. 2023), as seen in Figure 15 and 16. PoT applies different prompts for all datasets (Chen et al. 2023), they can found in Figures 17, 18, 19, 20 and 21. For Multi-Arith dataset, the original implementation only offered a Zero-Shot Approach, which we applied.